OpenStudy (anonymous):

Please someone help me. I'm stuck on this problem!!!Ricardo Gutierrez has a box of coins that he uses when playing poker with friends. The box currently contains 46 coins, consisting of pennies, dimes, and quarters. The number of pennies is equal to the number of dimes, and the total value is $5.26. How many of each denomination of coin does he have? The number of pennies, which is also equal to the number of dimes is... The number of quarters is...

OpenStudy (anonymous):

p + d + q = 46 p*1 + d*10 + q*25 = 526 In both equations sub p=d, so you'll have 2 eqns in p and q Solve for p and q, and get d

OpenStudy (anonymous):

So you would multiply 46 by each one

OpenStudy (anonymous):

p + d + q = 46 2p + q = 46 p*1 + d*10 + q*25 = 526 11p+25q=526 So now you have 2 eqns and 2 variables. Can you solve that.

OpenStudy (anonymous):

I very confused on this problem. Do I solve both equations on the bottom or stop.

OpenStudy (anonymous):

Could you show me the steps

OpenStudy (anonymous):

Just solve these two and find p and q 2p + q = 46 11p+25q=526

OpenStudy (anonymous):

Use a calculator if you have one that can do equations

OpenStudy (anonymous):

I don't have one

OpenStudy (anonymous):

p=16, q = 14 d = p = 16